OpenStudy (anonymous):

Which of the following is NOT in the solution set of 3d – 4 ≥ –19?

OpenStudy (anonymous):

\[3d-4\ge-19\iff3d \ge-15\iff d \ge-5\]

OpenStudy (texaschic101):

zairhenrique already gave you the answer ....any number that is LESS then -5 (a number that is to the left of -5, is NOT a solution. (EX. -6,-7,-8,-9, etc.)